Fermi National Laboratory


Accelerator Update

Wednesday June 2
The day shift began with Operations monitoring store 3552, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:15 AM, a Pbar Debuncher vacuum pump tripped off. It reset okay, but an expert looked at it and determined that there was a problem in the tunnel. He disconnected the control cable and will investigate further when an access becomes possible.

Tevatron (TeV) experts began an end-of-store study at 9 AM.

Operations established beam to the Meson MTest experiment at 10:16 AM.

Operations terminated store 3552 at 10:04 AM. Cryo system technicians were informed that they could begin their work on the TeV sector D3 wet engine. They completed the job a bit before noon.

TeV experts began a bump study at 10:22 AM.

The TeV studies ended and D-zero experimenters made a forty-five minute access into their collision hall at noon.

Operations put the TeV into shot setup at 1:51 PM. All Switchyard beam gets disabled during shot setups.

Operations established store 3554 at 3:57 PM with an initial luminosity of 70.0E30.

Pbar experts began a flying wire study in the Accumulator at 4:15 PM and ended the study at 4:52 PM. Pbar resumed stacking a couple of minutes later.

Operation put the Recycler into shot setup for stashing antiprotons at 7:26 PM. By 8:41 PM the Recycler had taken 34mA of antiprotons and Pbar resumed stacking.

At 8:55 PM, Operations reported having trouble getting beam out of the Main Injector (MI) for the fixed target users. Operators reset the QXR at 9 PM and this eventually resolved the problem. Operations reestablished beam to MTest.

At 10:28 PM, Operations reported that they lost the MI power supply permit. Everything cleared except for MI-60. The Duty Electrician investigated and discovered a 13.8kV Bus breaker trip for feeder 71. He called in high voltage experts.

Thursday June 3
The midnight shift began with operations monitoring store 3554 and the Recycler stash, and with MiniBooNE taking beam.

The high voltage expert reported at 12:09 AM that he had closed feeder 71 and was turning on MI. The experts couldn't find a feeder problem. They believed that something in MI-60 had brought the feeder down.

Stacking resumed at 12:12 AM.

At 11:22 AM, a Booster vacuum pump tripped off and caused two RF stations (BRF7 & 8) to trip. Everything reset okay.

Pbar resumed stacking at 12:46 PM, but very little beam was making it to the stacktail. Operators and an expert investigated, found the problem, and had Pbar stacking by 3:20 PM.

Pbar experts began a study period at 5:15 PM. Pbar continued stacking, but a reduced rate.

Pbar experts switched to a reverse proton study at 9 PM. They completed this study at 11:48 PM.

Friday June 4
The midnight shift began with Operations monitoring store 3554 and the stash, and with MiniBooNE taking beam.

Pbar resumed normal stacking at 12:02 AM.

At 3:49 AM, Operations reported that they had lost the Pbar stack due to a quad power supply (A:LQ) ground fault, caused by a LCW leak. Although this was a very unfortunate occurrence that stack had lasted for 69 days, 10 hours, and 8 minutes. There had been a stack in the Accumulator since March 26th, 2004. This broke the last record for stack longevity made back in December of 1987.

Operations prepared Pbar for an access at 5:31 AM, to occur while techs repaired the A:LQ power supply leak. The access began at 6 AM. The techs completed their repairs at 6:37 AM.

Plans for Today
The plans for today include terminating the store around 11 AM, and then allowing TeV experts to conduct six hours of studies. Operations will resume normal stacks and stores for the weekend. Recycler will stash antiprotons after the next store gets established later this evening.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/4/2004   email Fermilab